OpenStudy (ria23):

Verify the identity. cot(θ - pi/2) = -tan θ

OpenStudy (aum):

\(\cot(\theta - \pi/2) = -\cot(\pi/2 - \theta)\).  From the diagram: \(-\cot(\pi/2 - \theta) = -\tan(\theta) \)